OpenStudy (anonymous):

the sum of two numbers is seven. the sum of the squares of the two numbers is twenty nine. find the two numbers

OpenStudy (anonymous):

Well it is 2 and 5 cuz 2+5= 7 and 4+25= 29 But I guess they want you to do it this way: x+y= 7 solve for x here u get x= y-7 then plug here x^2 + y^2 = 29 u get (y-7)^2 + y^2 = 29 solve for y : Binomial theorem y^2 - 14y + 49 + y^2 = 29 ...etc